In still water , Larry sails his boat at an average of 30km/h. One day, he spent as much time sailing 104km down river as he did sailing 91km back up the river. What was the rate of the river current ?
Answer by stanbon(75887)

In still water , Larry sails his boat at an average of 30km/h. One day, he spent as much time sailing 104km down river as he did sailing 91km back up the river. What was the rate of the river current ?
-----------------
Let current rate be "c"
-----------------
Down River DATA:
Distance = 104 km ; Rate = (30+c) km/h ; Time = 104/(30+c) hrs
-------------------
Up River DATA:
Distance = 91 km ; Rate = (30-c) km/h ; Time = 91/(30-c) hrs
--------------------
EQUATION:
time up = time down
91/(30-c) = 104/(30+c)
104(30-c) = 91(30+c)
3120-104c = 2730+91c
195c = 390
c = 2 km/h (speed of the current)
==============
Cheers,
Stan H.